In a closed economy where the taxes are autonomous and marginal consumption propensity is X; how much increase on national income does generate an increase of Y billion public expenditures financed by an equal amount of tax increase?

(chose a value for x and y and make the calculation)
Answer #1

Let MPC = X = 0.8 (assumed)

Therefore, MPS = 1 – MPC = 1 – 0.8 = 0.2

Let government or public expenditure = Y = $2,000 billion

Hence,

Multiplier = 1/MPS = 1/0.2 = 5

National income = Y × Multiplier

                        = 2,000 × 5

                        = $10,000 billion (Answer)